OpenStudy (anonymous):

I am helping my nephew with his algebra. I'm pretty sure this is right just I want someone to double check it. Thanks.. x^4+x^2=12 x^4+x^2-12=0 (x^2+4)(x^2-3) x=+/-sqrt. of 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is true. you can use other variable to comfort (a^2+a-12)=0 (a+4)(a-3)=0 replace a=x^2 and ...

OpenStudy (anonymous):

@PhatDad

OpenStudy (anonymous):

thanks for the responses. I like the substitution, maybe it will help him see it.